OpenStudy (anonymous):

Find the sum of the Series: Sigma (-1)^(n-1) (3/4^n) from n=1 to infinity

OpenStudy (blockcolder):

You mean this: \[\sum_{n=1}^{\infty}(-1)^{n-1}\frac{3}{4^n}\]

OpenStudy (anonymous):

Yes

OpenStudy (blockcolder):

This is an alternating geometric series with r=-1/4 and a=3/4. Thus, the sum is \[\frac{\frac{3}{4}}{1+\frac{1}{4}}=\frac{3}{5}\].

OpenStudy (campbell_st):

the common ratio is -1/4 and 1st term a = 3/4 then \[s _{\infty} = a/(1-r)\]

OpenStudy (anonymous):

Thank you!

OpenStudy (blockcolder):

No problem. :D